### v1

# QCD Phase Diagram and Astrophysical Implications

Abstract
I make a brief review about the QCD phases and the equation of state inferred from the neutron star data. Along the temperature axis at low baryon density, the QCD phase transition is a smooth crossover, and it is a natural extension of our imagination to postulate a similar crossover along the density axis at low temperature. Even without phase transitions, the inferred thermodynamic properties of neutron star matter turn out to be highly nontrivial already at twice of the nuclear saturation density. I also give some discussions about the substantiation of quark matter by means of the gravitational wave signals including the multi-messenger prospect.
1 Important lessons we have learned from the heavy-ion collision
It is well-known that quantum chromodynamics (QCD) has two extreme limits that exhibit global symmetries to characterize the QCD phases. These symmetries are only approximate in the realistic case with physical parameters. In the limit of zero quark mass, chiral symmetry becomes exact, which is spontaneously broken at low energy as
$SU(N_{\mathrm{f}})_{\mathrm{L}}\times SU(N_{\mathrm{f}})_{\mathrm{R}}~{}\to~{}SU(N_{\mathrm{f}})_{\mathrm{V}}\,,$ (1)
where $N_{\mathrm{f}}$ is the number of quark flavors. In practice, $N_{\mathrm{f}}$ varies according to the typical energy scale of relevant processes as compared to the quark masses. For the environment in the neutron star, the quark chemical potential is less than $1\,\mathrm{GeV}$, so the active quark flavors are limited to the up, the down, and the strange. This corresponds to an intermediate situation between $N_{\mathrm{f}}=2$ and $N_{\mathrm{f}}=3$, which is often referred to as $N_{\mathrm{f}}=2+1$. In the opposite limit of infinite quark mass or the quenched limit of $N_{\mathrm{f}}=0$, center symmetry becomes exact, which is realized in the low-temperature phase and spontaneously broken in the high-temperature phase as
$Z_{N_{\mathrm{c}}}~{}\to~{}\mathbb{1}$ (2)
with the color number $N_{\mathrm{c}}$ that is $N_{\mathrm{c}}=3$ in QCD but could be artificially adjusted for theoretical purposes. If the realization of the global symmetries differs as physical parameters change, there must be at least one phase transition at some critical values of the parameters; For the temperature as the physical parameter, the critical value is denoted by $T_{\mathrm{c}}$. As a function of the quark mass $m_{\mathrm{q}}$, two phase transitions must be found at $m_{\mathrm{q}}=0$ and $m_{\mathrm{q}}=\infty$. Unfortunately, however, the above global symmetries cannot coexist for any $m_{\mathrm{q}}\neq 0$. The question is which phase transition is closer to the physical point.
This question has been intensely investigated in the context of the search for the quark-gluon plasma (QGP) at high temperature and low baryon density. Here, I only summarize what we know about the natures of the QCD phase transitions without going into phenomenological details about how they have been concluded.
— The chiral phase transition has not been observed.
The masses of the up and the down quarks are two orders of magnitude smaller than the typical QCD scale, i.e., $\Lambda_{\mathrm{QCD}}\simeq 0.2\,\mathrm{GeV}$. Therefore, it would be conceivable that the physical point may be not so far from the chiral limit of $m_{\mathrm{q}}\to 0$, and if so, some remnant phenomena associated with chiral symmetry breaking (1) should be expected. Strangely, however, no criticality has been confirmed by the experimental data from the heavy-ion collision. For this missing link between chiral symmetry and the heavy-ion collision data, there are three (and probably more) explanations.
The first one is that the critical region is too small. Although $m_{\mathrm{q}}\ll\Lambda_{\mathrm{QCD}}$ is certainly the case, the physical pion mass is comparable with the QCD scale, i.e., $m_{\pi}\simeq\Lambda_{\mathrm{QCD}}$, and the soft mode cannot be sufficiently soft around the pseudo-critical temperature. It should be noted that the pseudo-scalar channel is the lightest if the density is zero and the QCD inequality is applied (Nussinov and Lampert, 2002). The second explanation is that chiral symmetry affects hadron spectra, but does not drastically change the physical degrees of freedom, which implies that the bulk thermodynamics is not sensitive to the chiral phase transition. To observe consequences from the chiral phase transition, thus, some excitation on top of the bulk thermodynamics such as fluctuations and transport properties must be measured, but such quantities are rather fragile and easily diluted in the average over the whole dynamical evolution. The third problem is that good probes for chiral symmetry are not quite established, in my opinion. I know that people have discussed the medium modification of the vector mesons quantified by the dilepton measurements. Theoretically speaking, it is correct to say that, once the vector meson spectral function is somehow obtained, the expectation values of operators are constrained by the QCD sum rule. Even if the vector meson mass decreases, however, it does not uniquely lead to the decreasing behavior of the chiral order parameter; in other words, the Brown-Rho scaling is a convenient hypothesis but not a theorem. For a counter-example, see (Klingl et al., 1997). The vector meson may exhibit not the mass shift but only the width broadening, and then the connection to chiral symmetry is more subtle.
— The deconfinement phase transition has not been observed.
Then, one may want to think that the transition associated with the symmetry change (2) should have been probed, but it is not the case, either. In the absence of dynamical quarks, the pure gluonic theory has a first-order phase transition for $N_{\mathrm{c}}=3$, but dynamical quarks smear center symmetry of $Z_{N_{\mathrm{c}}}$. As a result, there is no sharp phase transition of deconfinement at all, while the thermodynamic quantities such as the energy density grow up within a narrow range of the temperature. In the past, some people were serious about the phenomenological implication from the symmetry breaking (2), that is, the dynamics of the QCD center domain walls associated with the discrete symmetry breaking in the early universe and the primordial magnetic field. Nowadays, however, it is a well-established fact that the deconfinement is a smooth phenomenon and there is no hope to salvage any hint about the center domain walls associated with deconfinement.
The absence of any sharp phase transition of deconfinement also implies a radical conjecture. Suppose that QCD has strict color confinement at zero temperature, then, what should be the probability to find quark excitation at low temperature, say, $\sim 0.01\,\mathrm{GeV}$? Is it strictly zero as literally imposed by color confinement or is it exponentially suppressed and yet nonzero? When I was a young postdoc, I asked this question to a renowned physicist and the answer I got was quite disappointing. I can never forget his answer — I don’t trust the finite-temperature formalism, that is why I wrote papers on the system in finite volume box, but not a single paper at finite temperature… At that time, I was young enough to stave off pessimism, and I kept my face up to write finite-temperature papers. Anyway, in spite of long-standing efforts, the order parameter for deconfinement has not been discovered except for the pure gluonic case. More precisely speaking, only the quark deconfinement order parameter (the Polyakov loop in the fundamental representation) is established and even in the pure gluonic theory the gluon deconfinement order parameter is not known or it does not exist; the Polyakov loop in the adjoint representation always takes a nonzero value. It would be not a bizarre idea to interpret the absence of the order parameter as the absence of phase transitional changes at all. Then, as soon as a finite temperature is introduced, the thermal excitation of colored objects may not be prohibited stringently. If this statement makes sense, the definition of the QGP would be obscure.
– The crossover is boring, but it is the real, and the most difficult to understand:
Theoreticians tend to idealize the system taking the limit of either $m_{\mathrm{q}}=0$ or $m_{\mathrm{q}}=\infty$. Critical phenomena are characterized by model independent analyses in theory and the experimental challenge to probe them is very appealing. However, the physical quark mass is not close to these limiting values, and the change from hadronic matter to the QGP takes place only smoothly if the baryon density is small. This smooth change is commonly referred to as the “crossover” in the QCD community. This usage of the terminology is analogous to that of the BEC-BCS crossover.
The scenario of no phase transition may sound like a boring reality, and indeed, this is the most boring in the phenomenological sense. There is no critical enhancement of fluctuations at the second-order phase transition, and there is no bubble nucleation at the first-order phase transition. There is no clear signature for the realization of the QGP, unfortunately.
Figure 1: Schematic picture of the crossover from hadronic matter to the QGP. The figure is adapted from (Fukushima, 2014).
Even though there is no exact way to define the QGP in quantum field theory, one could reasonably declare the creation of the QGP if any hadronic calculations cannot reproduce observed properties and perturbative QCD (pQCD) results can be consistent with them. The idea is visualized in Fig. 1. Actually, I emphasized this picture of the crossover in the proceedings contribution to Quark Matter 2014 (Fukushima, 2014). In the low-temperature phase, the hadron resonance gas (HRG) model agrees with the thermodynamic quantities measured in the lattice-QCD simulation, but the pressure from the HRG model eventually blows up. From the high-temperature side, the resummed perturbation theory of QCD works very well to explain the lattice-QCD thermodynamics up to some temperature, but the uncertainty is getting larger at lower temperatures. In this way, one can distinguish the hadronic phase well approximated by the HRG model and the QGP well described by the pQCD-based calculation.
One may wonder how these distinct regimes could be smoothly connected. In fact, that is the question. There are several preceding attempts along these lines. For example, the validity region of the HRG model can be extended toward higher temperature by the inclusion of the excluded volume effects (Andronic et al., 2012). Also, the resummed pQCD results near the pseudo-critical temperature may be ameliorated by the inclusion of the random gauge background modeled by the Polyakov loop averaging (Vuorinen and Yaffe, 2006). For some extension in terms of the semi-QGP, see (Hidaka and Pisarski, 2008). Then, it is likely that the improved hadronic and pQCD results may have the overlapping region of validity, in which one may say that matter contents are both hadronic and partonic (quarks/gluons). I must judiciously emphasize that I am not talking about a mixture of hadrons and partons. The identical physical system near the pseudo-critical temperature is equivalently regarded as either hadronic matter or partonic matter, and this concept is nothing but the duality in physics.
Although it is boring in phenomenology, the crossover is the most difficult to understand from the microscopic mechanism. In the duality region of the temperature, two descriptions in terms of hadrons and partons marginally work, and in turn, they nearly break down. Thus, besides the numerical simulations, there is no reliable theoretical device to investigate the matter properties in this intermediate region. Nevertheless, if one is interested in the thermodynamics only without caring about microscopic ingredients, one can construct it by a smooth interpolation between the HRG and the pQCD branches, as depicted in Fig. 1.
2 Not the best but the better among worse at high baryon density
We live on the first-order phase transition. Nuclei are self-bound fermionic systems. If the baryon chemical potential, $\mu_{\mathrm{B}}$, is too small, no baryonic matter can exist at zero temperature. With increasing $\mu_{\mathrm{B}}$ in isospin symmetric matter, a first-order phase transition occurs and the baryon density jumps from zero to the saturation density. This is a phase transition from the vacuum to a liquid of nuclear matter. Of course, we cannot live in a liquid of nuclear matter at the saturation density. So, our world lies exactly at the onset, $\mu_{\mathrm{B}}=m_{N}-B$, where $m_{N}$ is the lightest baryon (nucleon) mass and $B$ is the binding energy $\sim 0.016\,\mathrm{GeV}$ per nucleon.
The first-order phase transition is expected whenever such self-bound fermionic systems emerge. Therefore, if quark droplets are stable (or meta-stable), which hints the existence of the quark star, a first-order phase transition may be found at higher densities than the nuclear saturation density. This fascinating possibility cannot be ruled out so far, and the quark star hunt is one of the most exciting activities in nuclear and astro physics.
Figure 2: Crossover vs. first-order phase transition scenarios at high baryon density on the QCD phase diagram. The figure is adapted from Fujimoto et al. (2023).
Now, let us recall the lessons from the heavy-ion collision. In the past, many intriguing possibilities were considered, but after all, the most boring scenario turned out to be the case. Based on the wisdom we earned from high-temperature QCD matter, I would say, it should not be too conservative even if I postulate the least structured phase diagram, namely, only the crossover from hadronic matter to quark matter at high densities; see Fig. 2 for scenarios of the crossover and the first-order phase transition on the phase diagram. Because the duality plays such an essential role for deconfinement at high temperature, we must ask to ourselves; why not at high baryon density? Such an assumption of the least structured phase diagram may sound like the well-known principle of Occam’s razor, but I must say, the duality of deconfinement is a concrete physical concept and the crossover should conceal rich physics behind it.
It is still an attractive argument that quark matter might energetically be favored after a first-order phase transition, which is a stereotype scenario based on the bag model that provides us with an intuitive insight into the deconfining mechanism. However, again, we must ask to ourselves; if the bag model picture were validated at high density, why not at high temperature? It is already the fact that the QCD phase transition at high temperature and low baryon density is a smooth crossover and the duality connects two regimes in terms of different physical degrees of freedom. From this modern view of deconfinement, I believe that the crossover scenario at high density should be rather natural than others that require too specific assumptions, whereas the crossover scenario may not hit the exact ground truth. I remember that when I was a student at junior high school, an art teacher told us how to make a dessin, that is a rough sketch of drawing, of a human body for example. Then, first, one should be careful of the whole body and the balance of volumes and shapes of the body parts. It is not recommendable to draw full details of one arm only before placing the shoulder, the head, etc. It is sometimes perplexing to me that some people stick to minutiae such as an undenied possibility of weak first-order phase transitions before taking a whole perspective of the QCD phase diagram. Probably, a fine dessin of hands and fingers inveigles some theoreticians.
3 Likely EoSs supporting the crossover but the first-order phase transitions popping up
I am not so negative against the scenario of the first-order phase transition. I am simply saying that the crossover scenario is a good starting point to draw the first approximation for the equation of state (EoS), and then one can engrave this base with more structures. To make my point clear, let us take a close look at the inferred EoSs from the machine-learning analysis. We have invested resources to establish a method to solve the inverse problem from the neutron star data on the radius-mass ($R$-$M$) plane to the EoS candidate. The technique was first developed by means of the Bayesian analysis. It is extremely nontrivial in this method of the Bayesian analysis how to evaluate the conditioned probability to find the observation data for the given EoS, that is, the likelihood. Once the EoS is specified, one line of the combination of $(R,M)$ of the neutron star is fixed. Then, suppose that this $(R,M)$ line is the true answer, how can we quantify the probability to obtain the observation data with error bars? Actually, the distribution of the observation data may be significantly different from the Gaussian, and there is no simple machinery to quantify the conditioned probability. Therefore, some sort of the Monte-Carlo simulation is indispensable, and the entire procedures are time consuming.
Another machine-learning technique that utilizes the neural network demonstrates a good performance to predict an EoS candidate in response to a given observation dataset. AI allergic people say that the machine-learning method is a black box and the Bayesian analysis is superior. But we should keep in mind that the Bayesian analysis is one of the useful machine-learning methods, and we can choose the most convenient method for the problem we want to solve. Our final objective is to quantify the posterior distribution of the EoS, and we can achieve this goal using the neural network approach as follows. In the supervised learning, the model is trained by the training data, and the trained model predicts the output EoS for the input data. We can prepare independent trained models with independent training data. If the output is not well constrained, different models should lead to different predictions for the same input, and vice versa. We have conducted this numerical experiment (Fujimoto et al., 2021) and the results are shown in Fig. 3.
Figure 3: 100 EoSs inferred from 100 independently trained models. The input is the common real observation data of the neutron star radius and mass. The orange-colored plateau represents the approximately identified region of the first-order phase transition. The figure is adapted from (Fujimoto et al., 2021).
Among 100 outputs from independently trained models, in Fig. 3 we see that some flat regions occasionally appear. The orange lines indicate the regions with the slope, $c_{s}^{2}=dp/d\varepsilon$, less than 0.01, which can approximately be regarded as the first-order phase transition. If the EoS is parametrized by the piecewise polytrope with more segments, weaker first-order phase transitions can be more easily accommodated. The general tendency is that the lower density region may have a better chance to have the first-order phase transition, but the interpretation needs caution. For the piecewise polytrope, it is customary to make the partitions with equal interval in the logarithmic scale, and this means that the first-order phase transition in Fig. 3 gets stronger at higher density by construction of the piecewise polytrope. Therefore, naturally, the stronger first-order phase transitions at higher densities are less favored.
Here, a message I want to convey from Fig. 3 is not the relevance of the weak first-order phase transition but my view of the EoS construction; the overall behavior of the EoS is the averaged one which is very smooth and there is no characteristic point in the averaged global shape. Then, within the uncertainty band estimated by fluctuating outputs in Fig. 3, the EoS can gradually be updated by future data. Hereafter, I discuss the averaged EoS only but I never exclude the possibility of weak first-order phase transitions. Simply, the currently available data are not sufficient to resolve such ripples on top of the bulk EoS base.
Figure 4: Speed of sound squared as a function of the energy density in the unit of the energy at the saturation density, namely, $\varepsilon_{0}=0.15\,\mathrm{GeV}/\mathrm{fm}^{3}$. The dependence on the number of segments, $N$, is only minor. The figure is adapted from Fujimoto et al. (2024b).
The slope of the EoS is the speed of sound squared, i.e., $c_{s}^{2}=dp/d\varepsilon$, and this quantity approaches $1/3$ in the conformal limit. If there is no mass scale other than the chemical potential, the pressure is parameterized with a numerical coefficient as
$p=\nu\mu_{\mathrm{B}}^{4}\,,$ (3)
and then the energy density is $\varepsilon=\mu_{\mathrm{B}}(dp/d\mu_{\mathrm{B}})-p=3\nu\mu_{\mathrm{B}}^{4}$. Thus, one can immediately conclude $c_{s}^{2}=1/3$ in this case. The latest results of the speed of sound from the machine-learning inference are presented in Fig. 4. Although the error bar (the $1\sigma$ band) is not small, it is fair to say that a peak may stand around $\varepsilon\simeq 4\varepsilon_{0}\simeq 0.6\,\mathrm{GeV}/\mathrm{fm}^{3}$. I am a bit prudent about the statement. I was once sloppy enough to give a talk and use a word, peak, by showing Fig. 4. Then, an experimental colleague in the audience immediately reacted to refute my careless statement and said that no peak was visible. It was experimentally a correct remark, but as a theoretician, I would still insist that this should be a peak even though the peak width might be broad. Since QCD is an asymptotically free theory, the speed of sound squared must go to the conformal value, $1/3$, eventually in the high density limit.
It is actually a nontrivial question whether the slightly decreasing behavior for $\varepsilon/\varepsilon_{0}>4$ is physical or artificial. In the older analysis in (Fujimoto et al., 2021), we were inclined to consider that the decrease was attributed to the machine-learning artifact. If there is no constraint from the data, the output should be inherited from the distribution of the speed of sound squared in the training data, that is, $c_{s}^{2}\sim 0.5$. However, in the follow-up analysis in (Fujimoto et al., 2024b), we have confirmed that the decreasing behavior even for $\varepsilon/\varepsilon_{0}>4$ is well constrained by the observation data. This sounds neat; however, this is a very puzzling assertion. From the astrophysical constraint, the soft EoS can be excluded by the presence of massive neutron stars, but the stiffer EoSs are always allowed as long as it does not violate the causality bound. Physically speaking, there is no reason why $c_{s}^{2}$ should decrease for $\varepsilon/\varepsilon_{0}>4$. Among the authors of (Fujimoto et al., 2024b), we had a lot of discussions, and our hypothetical conclusion is that the machine-learning model might give an additional meaning to the “absence” of further massive neutron stars. From the discovery of something, usually, people can impose a constraint so that this something can exist. But, from the non-discovery, we have no idea whether something does exist but it is not discovered yet, or this something does not exist at all. Therefore, we usually do not care about the non-discovery, but the machine-learning model does. Actually, it is not an insane idea to take the non-discovery such actively. In science, some principles, desirably simpler principles, are assumed and they are accepted unless some counter-example is found. In the same way, we may well assume that the non-discovery signifies the non-existence. Under this assumption, the heaviest neutron star is to be identified as the upper limit and the EoS cannot be stiff to exceed this upper limit. One may say that this is just an artifact from the machine-learning inference, but I myself was thrilled about the potential of the machine-learning approach that may break out of our shell of thinking.
4 Trace anomaly approaching zero means what?
Once the EoS is somehow inferred, various combinations of $p$ and $\varepsilon$ can be considered, and the speed of sound squared is only one example. Another interesting quantity is the conformality measure defined by
$\Delta=\frac{1}{3}-\frac{p}{\varepsilon}=\frac{\varepsilon-3p}{3\varepsilon}\,.$ (4)
The numerator, $\varepsilon-3p$, is the trace of the energy-momentum tensor. The energy-momentum tensor and the dilatation current can be improved so that the divergence of the dilatation current can be expressed by the trace of the energy-momentum tensor. Therefore, if the dilatation current is conserved due to conformal symmetry, the trace of the energy-momentum tensor is vanishing. In the classical level QCD is an almost scale free theory except for the quark masses, but the running coupling involves an additional scale, $\Lambda_{\mathrm{QCD}}$, from the quantum effects. Then, the trace of the energy-momentum tensor is no longer vanishing but it is given by a combination of the gluon condensate and the $\beta$ function that characterizes the running coupling.
Here again, let me explain what is known about the trace anomaly at high temperature.
— The trace anomaly in the matter part is peaked around the critical temperature.
In lattice QCD, the dimensionless quantity, $(\varepsilon-3p)/T^{4}$, is often called the interaction measure. In the pure gluonic theory (Boyd et al., 1996) as well as QCD, a peak in the interaction measure stands near the critical temperature. It would be misleading to say that the trace anomaly is enhanced near the critical point, though. The fact is rather opposite. The vacuum part of the gluon condensate is positive, and the gluon condensate melts around the critical temperature, which means that the negative matter part partially cancels the positive vacuum part. Then, for the temperatures higher than $T_{\mathrm{c}}$, the gluon condensate goes to negative (the trace anomaly is positive). Therefore, the correct interpretation is that the trace anomaly including the vacuum part is not enhanced but crossing the zero near the critical temperature.
At higher temperatures, one may think that the interaction measure should be suppressed as $T^{-4}$, but this is not really the case. It was recognized that the interaction measure scales as $T^{-2}$ at high temperature, which implies that the finite-temperature gluon condensate has a term behaving nonperturbatively like $\Lambda_{\mathrm{QCD}}^{2}T^{2}$ even at high temperature.
Figure 5: Conformality measure from the neutron star data (solid line) and the finite-temperature lattice-QCD data (dashed line). The horizontal axis is the energy in the unit of $\varepsilon_{0}=0.15\,\mathrm{GeV}/\mathrm{fm}^{3}$.
The pressure, $p$, as a function of the temperature, $T$, is parametrized in (Bazavov et al., 2014) based on the first-principles calculation, and it is easy to compute the conformality measure (4). In the QCD case with dynamical fermions, the peak is smeared as shown by the dashed line in Fig. 5, while a sharp peak appears in the pure gluonic theory.
A surprise awaits to be found when we consider $\Delta$ using the neutron star data.
— The trace anomaly in the matter part rapidly decreases in the central cores of the neutron star.
The solid line in Fig. 5 shows $\Delta$ corresponding to the neutron star EoS in Fig. 4, which is parametrized in (Fujimoto et al., 2022). I would point out that there are (at least) two surprises. One is that the characteristic energy scales are surprisingly small and they are relevant to the realistic neutron star environment. We see that $\Delta$ rapidly decreases already around $\varepsilon/\varepsilon_{0}\gtrsim 2$. Then, around $\varepsilon/\varepsilon_{0}\sim 5$ which is reachable in the deepest core of the heavy neutron star, the conformality measure is strongly suppressed and, it could even go to negative, which is the second surprise.
At finite density and low temperature, the conformality measure starts from $\sim 1/3$ in the low-density side, and this can be understood from $\varepsilon\gg p$ in the nonrelativistic regime. In other words, along the finite density direction, the nucleon mass significantly breaks scale invariance and $\Delta$ is naturally enhanced there. Unlike the finite temperature case, it is subtle to interpret the behavior of this matter part in terms of the gluon condensate. One may say that the trace anomaly has not only the gluon condensate but also the chiral condensate, but the latter disappears in the massless quark limit. I believe that the qualitative and even the quantitative nature of $\Delta$ as shown in Fig. 5 would hardly be changed in the massless limit. Naïvely in the QCD language, together with the vacuum part, the positive gluon condensate seems to increase further as the density grows up. This is unnatural at first glance, but this seemingly unnatural behavior could be justified if some forms of new condensates develop with increasing density.
The necessity of additional condensates is strongly motivated by the following argument as well. If $\Delta$ goes to negative, it is straightforward to see,
$\varepsilon-3p=\mu_{\mathrm{B}}^{5}\frac{d}{d\mu_{\mathrm{B}}}\frac{p}{\mu_{\mathrm{B}}^{4}}<0\,,$ (5)
which means $d\nu/d\mu_{\mathrm{B}}<0$ where $\nu$ is the thermal degrees of freedom introduced in Eq. (3). This is an incredible statement; the thermal degrees of freedom decreases with increasing chemical potential. How can it be true? Interestingly enough, there are several known examples in which $\varepsilon-3p<0$ is realized. The most well-known system is the high isospin matter with pion condensation (Son and Stephanov, 2001). The chiral perturbation theory should work as long as the isopin chemical potential is not too large, and $\varepsilon-3p<0$ is predicted, which has been confirmed by the lattice simulation (Abbott et al., 2023). It is interesting to point out that the nonperturbative quadratic term such as $f_{\pi}^{2}\mu_{\mathrm{I}}^{2}\sim\Lambda_{\mathrm{QCD}}^{2}\mu_{\mathrm{I}}^{2}$ is indispensable to account for the enhanced behavior of the speed of sound (Chiba and Kojo, 2024) and also for the negative conformality measure. At high temperature and high density, respectively, the manifestation of $\Delta$ looks very different, but in both cases the nonperturbative quadratic term $\sim\Lambda_{\mathrm{QCD}}^{2}T^{2}$ or $\sim\Lambda_{\mathrm{QCD}}\mu^{2}$ should be required. The microscopic origin might be common.
In the neutron star matter, it is likely that $\Delta$ crosses zero and $\Delta<0$ is realized around $\varepsilon/\varepsilon_{0}\simeq 4$-$5$. The promising candidate for the condensate there is the color-superconducting gap. Because the coupling constant is too small, the effect of the color superconductor is only moderate at high density where the pQCD treatment is justified (Fukushima and Minato, 2024). For the consistent interpretation of the neutron star EoS, nonperturbative enhancement of some condensates must be taken into account, which is a challenging problem.
5 Scanning the phase transitions with the total mass of the binary neutron star system
The merger of the binary neutron stars is such an explosive event that the future gravitational wave measurement is expected to give us useful information about extremely dense matter even for $\varepsilon\gtrsim 5\varepsilon_{0}$. Also in the intermediate density region, $\varepsilon\lesssim 5\varepsilon_{0}$, the tidal deformability in the inspiral stage before the merger can constrain the EoS well, and the uncertainty band in Fig. 3 should become much narrower soon. This is actually why I tend to claim that the fine details of the EoS should be updated later when more experimental data will be available, which will happen in the near future.
Figure 6: Schematic illustration of the phase transition scan by the gravitational wave signals. The softening from nuclear matter (NM) to quark matter (QM) changes the life time of the remnant after the merger. If the EoS is stiff, the hypermassive (or supramassive) neutron star (HMNS) remains, while the soft QM EoS cannot support it and the gravitational collapse to the blackhole (BH) is expected.
It is the most ambitious attempt to utilize the gravitational wave signals for the purpose of the detection of the QCD phase transitions. For the moment, nobody knows the exact phase structure at high density, and my belief is the crossover from hadronic to quark matter as emphasized previously. As a matter of fact, the order of the phase transition itself is not so important in this context. If the pQCD estimate of the EoS is extrapolated toward the neutron star density, the pQCD-based EoS is very soft, and the speed of sound squared cannot be far from the conformal value. For the fate of the binary neutron star merger, it is essential that the EoS of dense matter after some phase transition or crossover is expected to be soft, and it is almost irrelevant how to transmute the matter with or without a discontinuous jump. The bottom line is the following: the QCD phase transition, which may be a first-order phase transition (Most et al., 2019; Weih et al., 2020) or a continuous crossover (Fujimoto et al., 2023), softens the EoS and the gravitational collapse to the blackhole is triggered by the onset of quark matter. The idea is schematically illustrated in Fig. 6. If the full shapes of the gravitational waves from the post-merger stage will be probed in the future, they would provide us with valuable information about the nature of the phase transitional change to quark matter. However, the nature is not so kind. After the merger, the system size is far smaller than the inspiral stage, and the amplitude of the gravitational waves is much reduced. The typical frequency range of the gravitational waves of our interest to investigate the QCD phase transition is around $\gtrsim 3\,\mathrm{kHz}$ and I am not very optimistic about the possibility to achieve the sensitivity of the detector in this frequency range. To confirm the QCD phase transition using the post-merger gravitational wave signals, we should be very lucky or we should complete the mission impossible to make a reliable and precise prediction or we should wait for decades till the detector sensitivity after several generations will be sufficiently improved.
Figure 7: Life time of the remnant after the merger as a function of the total mass of the binary neutron star system. CO represents the intermediate-density crossover scenario and PT represents the high-density first-order phase transition scenario. The figure is adapted from (Fujimoto et al., 2024a).
One workable breakthrough to overcome the sensitivity problem at high frequency band is the multi-messenger analysis. In the case of GW170817, an electromagnetic counterpart, AT2017gfo, identified as the kilonova has been observed. The extensive analysis of the kilonova gave us a lot of information about the r-process elements (Domoto et al., 2021). Here, I want to emphasize that the presence of the kilonova already imposes a condition onto the EoS candidate. The point is that the life time of the remnant after the merger should be long enough to acquire the disk mass in a way consistent with the estimated kilonova mass. Any phase transition scenario that cannot explain the kilonova mass must be strictly ruled out. If the luminosity of many observed kilonovae will be systematically investigated in the future, the phase transition scenarios can be discriminated even without looking at the post-merger signals. In fact, Fig. 7 shows the life time of the remnant after the merger as a function of the total mass of the binary neutron star system (Fujimoto et al., 2024a). CO and PT are two representative EoS scenarios. It is evident that the life time strongly depends on the EoS scenarios in a specific window of the total mass (labeled as “sweet spot” in Fig. 7). For example, according to Fig. 7, if a bright kilonova will be discovered for the merger with the total mass at $2.9M_{\odot}$, then the CO scenario must be denied. We note that, if the binary system is too light, the hypermassive neutron star lives long enough and the EoS differences are irrelevant. Also, if the binary system is too heavy, the remnant is immediately collapsed into the blackhole regardless of the EoS differences. The sweet spot is centered around twice of the typical neutron star mass $\sim 2.8M_{\odot}$, which was indeed the case for GW170817. In this way, the location and the strength of the phase transition including the crossover will be able to be constrained by the future observation in the total mass window of the sweet spot. In some events, the kilonova might be simply overlooked experimentally, but not always so. I am very optimistic about the systematic survey of the kilonova which will be able to render a verdict of the quark matter transition.
Acknowledgments
This proceedings contribution is based on my published papers as well as ongoing projects. The author thanks his collaborators of the published papers: Yuki Fujimoto, Kenta Hotokezaka, Syo Kamata, Koutarou Kyutoku,
Larry McLerran, Shuhei Minato, Koichi Murase, and
Michal Praszalowicz. He also thanks his collaborators of the ongoing projects: Len Brandes, Kei Iida, and Chengpeng Yu.